# Adv-BMT: Bidirectional Motion Transformer for Safety-Critical Traffic Scenario Generation

### v1

@@ Abstract
Scenario-based testing is essential for validating the performance of autonomous driving (AD) systems. However, such testing is limited by the scarcity of long-tailed, safety-critical scenarios in existing datasets collected in the real world. To tackle the data issue, we propose the Adv-BMT framework, which augments real-world scenarios with diverse and realistic adversarial interactions. The core component of Adv-BMT is a bidirectional motion transformer (BMT) model to perform inverse traffic motion predictions, which takes agent information in the last time step of the scenario as input, and reconstruct the traffic in the inverse of chronological order until the initial time step. The Adv-BMT framework is a two-staged pipeline: it first conducts adversarial initializations and then inverse motion predictions. Different from previous work, we do not need any collision data for pretraining, and are able to generate realistic and diverse collision interactions. Our experimental results validate the quality of generated collision scenarios by Adv-BMT: training in our augmented dataset would reduce episode collision rates by 20% compared to previous work. The code will be made available at https://metadriverse.github.io/adv-bmt/.
@@ 1 Introduction
In recent years, autonomous driving (AD) agents have achieved unprecedented performance in urban environments dosovitskiy2017carlaopenurbandriving ; li2022metadrivecomposingdiversedriving ; wu2024metaurbanembodiedaisimulation . However, handling corner traffic situations, especially collision scenarios, remains a major challenge. One cause is that safety-critical scenarios are missing from real-world driving datasets due to high costs and risks of data collections. Without enough collision training data, it is hard for the autonomous driving (AD) planners and prediction models to learn safe driving in challenging and risky scenarios. This motivates the need for simulating collision traffic that emulates real-world accidents. According to rempe2022generatingusefulaccidentpronedriving , useful safety-critical scenarios should be realistic and diverse, as well as challenging. Previous works rempe2022generatingusefulaccidentpronedriving ; zhang2023catclosedloopadversarialtraining ; stoler2025sealsafeautonomousdriving ; ransiek2024goosegoalconditionedreinforcementlearning leveraged learned real-world traffic priors to constraint adversarial trajectories, and optimized predictions on collision-encouraging objectives. However, from our observations and evaluations, generated behaviors from these methods have a lack of diversity and a lower collision success rates (attacking efficiency).
To improve upon previous work, we propose the Adv-BMT framework for realistic and diverse collision augmentations from real-world driving data. Adv-BMT samples various collision situations from the input driving log, and then reconstruct complete adversarial trajectories accordingly. This design enhances the diversity of predicted collision interactions, since now it augments different collision behaviors for each driving log. Besides, there is no need to include collision data in trainings. We further implement a rule-based rejection sampling mechanism to filter out candidate trajectories from unsatisfactory adversarial initializations, and maintain realistic collision interactions. In the core of Adv-BMT is the bidirectional motion transformer (BMT) model. Similar to existing multi-agent simulators and prediction models wu2024smartscalablemultiagentrealtime ; seff2023motionlmmultiagentmotionforecasting ; zhou2024behaviorgptsmartagentsimulation , BMT tokenizes continuous agent trajectories into discrete actions, and performs next-token predictions in an auto-regressive manner. Different from existing models, BMT learns to predict both future and history motion for all agents, conditioning on one-step current states. This is achieved by processing two sets of motion tokens from the input scenario, with one set of tokens used for forward prediction (future motion) and one for reverse prediction (history motion).
@@
As a summary, Adv-BMT is a two-staged pipeline: first, it initializes diverse collision frames between an new adversary agent (ADV) and ego vehicle; then, it reconstructs the adversarial trajectories via BMT’s reverse predictions. In the output, ADV is added to the original scenario, maintaining realistic interactions with surrounding traffic. An overview of our framework is illustrated in Fig. 1. We summarize our contributions as follows: (1) We introduce the bidirectional motion transformer with temporally reversible motion tokenizations; (2) We develop Adv-BMT for realistic and diverse safety-critical traffic simulations; (3) We leverage Adv-BMT in a closed-loop setting to dynamically create challenging environments for reinforcement learning agents.
@@ 2 Related Work
Driving Motion Predictions and Simulations
Recent work have utilized auto-regressive sequence modeling on discretized motion representations via transformer-based models. In previous works, Trajeglish philion2024trajeglishtrafficmodelingnexttoken discretizes motion representations using a k-disk based tokenization scheme to represent position and angle differences for relative movements; but the interaction performance is sensitive to agent ordering, and agents decoded first might not be as reactive to the rest of the predicted agents. MTR++ shi2024mtrmultiagentmotionprediction directly represents motion on continuous space in Gaussian mixture distributions. The design is efficient for multi-modal generations but prediction diversities across modes rely on the learned intention queries. MotionLM seff2023motionlmmultiagentmotionforecasting and SMART wu2024smartscalablemultiagentrealtime model motion actions that represents trajectory deltas on decoder-only transformers. BehaviorGPT zhou2024behaviorgptsmartagentsimulation performs next-patch predictions with future motion chunks, instead of single-step predictions, and achieves higher inference efficiency and reduces compounding errors.
Safety-Critical Scenario Generations
@@
Previous works leverage learned traffic prior to generate realistic safety-critical interactions. STRIVE rempe2022generatingusefulaccidentpronedriving enhances diversity through modeling traffic a latent space using variational autoencoders but requires computationally expensive per-scenario optimizations. CAT zhang2023catclosedloopadversarialtraining utilizes the augmentation framework into MetaDrive li2022metadrivecomposingdiversedriving simulator as a closed-loop reinforcement learning environment. Upon CAT, SEAL stoler2025sealsafeautonomousdriving uses a skill-based adversarial policy for human-like driving guided by collision-related objectives. MixSim Suo_2023_CVPR generates realistic reactions by exploring future goals and re-simulates the inferred trajectories using a trained policy. GOOSE ransiek2024goosegoalconditionedreinforcementlearning uses a goal-conditioned reinforcement learning approach, which parameterizes trajectories and goal constraints to generate scenario-level, safety-critical interactions. AdvSim wang2023advsimgeneratingsafetycriticalscenarios directly perturbs actor trajectories using a kinematic model and optimizes via a black-box adversarial loss. Recent work ding2020multimodalsafetycriticalscenariosgeneration uses a conditional normalizing flow to model the distribution of real-world safety-critical trajectories. SafeSim chang2024safesimsafetycriticalclosedlooptraffic proposes a diffusion model with a test-time collision-sensitive guidance loss to control the collision type and adversarial agent selections; however its generation quality relies on initial collections of collision data. CrashAgent li2025crashagentcrashscenariogeneration and LCTGen tan2023languageconditionedtrafficgeneration leverage free-form texts as inputs and extract embeddings to parameterize scene initializations and agent driving directions.
Figure 1: Overview of Adv-BMT framework. From the input scenario, Adv-BMT first samples diverse collision states between ego agent and a new adversarial agent, then reconstruct complete adversarial trajectories via BMT’s reverse motion predictions. A rule-based reject sampling mechanism is used for filtering unsatisfactory collision initializations.
3 Methodology
In this section, we formulate forward and reverse motion predictions. We then introduce bidirectional motion transformer (BMT) model. Last but not the least, we demonstrate details of Adv-BMT.
3.1 Problem Formulation
@@
Consider a traffic scenario with maximumly $N$ agents and time horizon $T$. The trajectory of agent $i$ is denoted by $\boldsymbol{\tau}^{i}=\{{\tau}^{i}_{0},{\tau}^{i}_{1},\dots,{\tau}^{i}_{T}\}$, where each state $\tau^{i}_{t}\in\mathbb{R}^{d}$ includes position, velocity, and heading at time $t$. We define a prediction direction indicator $D\in\{\text{Forward},\text{Reverse}\}$, for future or past motion prediction within $T$ time steps. For each scenario we construct a sequence of motion tokens:
$\mathbf{Z}^{i}=\{z^{i}_{1},\dots,z^{i}_{T}\},\quad z^{i}_{k}=\begin{cases}\phi(\tau^{i}_{k},\tau^{i}_{k-1}),&D=\text{Forward}\\ \phi(\tau^{i}_{T-k},\tau^{i}_{T-k+1}),&D=\text{Reverse}\end{cases}$ (1)
where $\phi(\cdot)$ denotes the motion tokenization function which we discuss in section 3.2. The goal is to model the distribution of agent motion tokens at the $k$-th step, conditioned on the predicted motion token sequence, current state, and scene context $\mathcal{M}$:
$P(z^{1}_{k},z^{2}_{k},\dots,z^{N}_{k}\mid\mathbf{z}^{1}_{1:k-1},\mathbf{z}^{2}_{1:k-1},\dots,\mathbf{z}^{N}_{1:k-1},\mathcal{M},D)$ (2)
where $\mathbf{Z}^{i}_{0:k-1}$ denotes history token sequence for agent $i$. The decoder autoregressively selects the next action token $z^{i}_{k+1}$ by attending to past tokens and scene token embedding $\mathcal{M}$. Each selected action in forward prediction maps to a position in future time-step, similar to a standard motion predictor; selected token sequence in reverse prediction can be reconstructed as history trajectories. Together, BMT is able to perform both future and past motion predictions as shown in Fig. 4.
Figure 2: BMT model architecture.
3.2 Bi-direction Motion Tokenizations
Token space
@@ 3.2 Bi-direction Motion Tokenizations
Control token simulation
Given a motion control token, BMT simulates it and reconstructs to a new agent state using midpoint integration:
$\text{Recons}(z,\tau,\Delta t)=\left(x+\bar{v}\cdot\cos(\bar{\theta})\cdot\Delta t,\;y+\bar{v}\cdot\sin(\bar{\theta})\cdot\Delta t,\;\theta+\omega\cdot\Delta t,\;v+a\cdot\Delta t\right)$ (3)
where $\tau=(x,y,\theta,v)$ denotes the current agent state with position $(x,y)$, heading $\theta$, and speed $v$; $z=(a,\omega)$ is the motion token containing acceleration $a$ and yaw rate $\omega$; $\Delta t$ is the time interval; $\bar{v}=v+\frac{1}{2}\cdot a\cdot\Delta t$ is the midpoint speed; and $\bar{\theta}=\theta+\frac{1}{2}\cdot\omega\cdot\Delta t$ is the midpoint heading.
Ground-truth tokens
For obtaining ground truth motion tokens during trainings, we simulate all $K^{2}$ candidate motion tokens starting from each agent state at time $t$, then select the token whose predicted next state $\hat{\tau}_{t+1}$ best matches the ground truth $\tau_{t+1}$. This process maps the continuous trajectory into a sequence of discrete motion tokens $\{z_{0},z_{1},\dots,z_{T}\}$ for each predicted agent. Forward and Reverse predictions share the same motion token space but differ in token interpretations. In reverse predictions, we invert the temporal order of the trajectory and treat the final state (e.g., collision initializations from Adv-BMT) as the initial; at each time step, we simulate each motion token (by inverting the time step, $\Delta t\to-\Delta t$), and find the token that best reconstructs the previous state $\tau_{t}$ from $\tau_{t+1}$. This generates a motion token sequence in reverse time direction $\{\tilde{z}_{0},\tilde{z}_{1},\dots,\tilde{z}_{T}\}$, where $\tilde{z}_{0}$ corresponds to the final state in forward time.
3.3 Token Matching with minimized Contour Error
For either reverse or forward predictions, we build motion tokens step-by-step in the order of prediction time, starting from the first prediction step and reconstructing till the last prediction step. With a 0.5-second time interval, a search is conducted from all candidate token vocabularies, and selects the best candidate that minimizes the distance between the ground-truth and reconstructed contours for the $i$-th candidate action for agent $n$:
$a_{n}^{*}=\arg\min_{i}\left(\mathcal{E}_{n}^{(i)}=\frac{1}{P}\sum_{p=1}^{P}\left\|\mathbf{c}_{n}^{(i,p)}-\hat{\mathbf{c}}_{n}^{(p)}\right\|_{2}\right)\quad\text{where }0\leq i<K^{2}$ (4)
where $\mathbf{c}_{n}^{(i,p)}\in\mathbb{R}^{2}$ and $\hat{\mathbf{c}}_{n}^{(p)}\in\mathbb{R}^{2}$ represent the $p$-th polygon corner of the reconstructed and ground-truth contours, each with totally $P$ contour points; $a_{n}^{*}$ is the index of the selected candidate token.
3.4 Bidirectional Motion Transformer (BMT)
The BMT architecture and training process is shown in Fig. 2, along with ground-truth token matching for forward predictions or reverse predictions.
Architecture
@@ 3.4 Bidirectional Motion Transformer (BMT)
BMT has a scene encoding component used to obtain embeddings for scenario contexts with separate embeddings for map polylines, traffic lights, and agent initial states. Then, we use Fourier-encoded edge features tancik2020fourierfeaturesletnetworks to represent the spatial and directional information between these encoded entities, which are then passed to the transformer encoder with self-attention layers for the relational embeddings.
@@
The prediction decoder predicts subsequent motion tokens in an autoregressive manner, with only initial agent information for the first frame, along with the scene embedding obtained from the Scene Encoder. The motion decoder incorporates self-attention over the initial agent token embeddings, and three relation computations separately: agent-to-agent (a2a), agent-to-time (a2t), and agent-to-scene (a2s), with each relation embedding then passed to its cross-attention layers. The output agent embeddings are concatenated and repeated a number of times. The output agent motion embeddings are mapped to the vocabulary of discretized motion tokens through MLPs.
Training
In training, BMT predicts the sequence of motion tokens and use the cross-entropy objective to match the joint action distribution of the observed behaviors in the dataset during training:
@@
$\max_{\pi_{\theta}}\;\mathbb{E}_{\mathcal{D}}\left[\sum_{t=1}^{T}\sum_{i}^{N}\log\pi_{\theta}\left(z^{i}_{t}\mid\mathbf{z}^{i}_{1:k-1},\mathcal{M},D\right)\right]$ (5)
During inference, BMT generate motion tokens autoregressively, and uses nucleus sampling (top-$p$ sampling) to enhance diversity and maintain plausibility. To prevent the exposure bias during auto-regressive inference, BMT rolls out upon chosen motion tokens instead of the ground-truth token sequence.
3.5 Adv-BMT
The overview of Adv-BMT is illustrated in Fig.1. The framework mainly consists of three steps: (1) it adds an adversarial agent (ADV) with a specified collision state into the current scenario; (2) it predicts reversely for the adversarial trajectory; (3) it performs rule-based checks and reject physically implausible ones.
Diverse Adversarial Initializations

### v2

@@ Abstract
Scenario-based testing is essential for validating the performance of autonomous driving (AD) systems. However, such testing is limited by the scarcity of long-tailed, safety-critical scenarios in existing datasets collected in the real world. To tackle the data issue, we propose the Adv-BMT framework, which augments real-world scenarios with diverse and realistic adversarial traffic interactions. The core component of Adv-BMT is a bidirectional motion transformer (BMT) model to perform inverse traffic motion predictions, which takes agent information in the last time step of the scenario as input, and reconstruct the traffic in the inverse of chronological order until the initial time step. The Adv-BMT framework is a two-staged pipeline: it first conducts adversarial initializations and then inverse motion predictions. Different from previous work, we do not need any collision data for pretraining, and are able to generate realistic and diverse collision interactions. Our experimental results validate the quality of generated collision scenarios by Adv-BMT: training in our augmented dataset would reduce episode collision rates by 20%. Demo and code are available at https://metadriverse.github.io/adv-bmt/.
@@ 1 Introduction
In recent years, autonomous driving (AD) agents have achieved unprecedented performance in simulations dosovitskiy2017carlaopenurbandriving ; li2022metadrivecomposingdiversedriving ; wu2024metaurbanembodiedaisimulation . However, handling corner traffic situations, especially collision scenarios, remains a major challenge. A major cause is that safety-critical scenarios are missing from real-world driving datasets due to high costs and risks of data collections. Without enough collision training data, it is hard for the autonomous driving (AD) planners and prediction models to learn safe driving in challenging and risky scenarios. This motivates the need for simulating different real-world accidents. To generate realistic collision trajectories, previous works rempe2022generatingusefulaccidentpronedriving ; zhang2023catclosedloopadversarialtraining ; stoler2025sealsafeautonomousdriving leveraged learned real-world traffic priors, and optimized predictions on collision-encouraging objectives. However, our evaluations of these baselines reveal that the generated behaviors are insufficiently diverse and yield a low collision generation rate.
Figure 1: Overview of Adv-BMT. The framework mainly consists of three steps: (1) it adds an adversarial agent (ADV) with a sampled collision state into the current scenario; (2) it predicts reversely for the adversarial trajectory; (3) it performs rule-based checks and rejects physically implausible ones.
We tackle this challenge by rethinking the motion prediction architecture itself. We introduce the Bidirectional Motion Transformer (BMT), a new model that learns to predict both future and history trajectories for all agents, conditioned on their current states. Similar to recent autoregressive traffic forecasting models wu2024smartscalablemultiagentrealtime ; seff2023motionlmmultiagentmotionforecasting ; zhou2024behaviorgptsmartagentsimulation , BMT tokenizes continuous trajectories into discrete control actions. Distinct from prior work, BMT employs a temporally reversible tokenization scheme that enables unified forward (future) and reverse (history) motion prediction within the same framework.
We utilize BMT model into our Adv-BMT framework for realistic and diverse collision traffic generations from real-world driving data. While existing works follow a standard paradigm, which selects a convenient neighbor agent and modifies the behavior to attack the ego agent, Adv-BMT inserts new agents (ADV) with diverse collision interactions, maintaining realistic interaction with other traffic agents. In short, Adv-BMT is a three-staged pipeline: first, it initializes diverse collision frames between a new adversary agent (ADV) and ego vehicle; then, it reconstructs the adversarial trajectories via BMT’s reverse predictions; finally, it conducts rule-based checks and rejects trajectories with physically implausible collision initializations. An overview of our framework is illustrated in Fig. 1. It is worth noting that that there is no collision data included in our model training.
Adv-BMT is designed with multiple generation modes to support varying levels of agent interactions and realism. By default, traffic agents follow their recorded trajectories to preserve consistency with the real-world driving log, while the adversarial agent is generated to interact within this fixed context. This enables targeted scenario editing while maintaining overall scene plausibility. To address limitations in interactivity, Adv-BMT also supports a closed-loop reverse prediction mode, in which all agents are jointly predicted to generate a fully reactive traffic scenario, without teacher-forcing any agent behaviors during predictions. Additionally, an optional forward refinement step allows traffic agents to respond to the newly introduced adversary, enabling a more interactive and dynamically consistent outcome. Together, these modes allow Adv-BMT to balance realism, controllability, and diversities for different use cases.
@@
We summarize our contributions as follows: (1) We introduce the bidirectional motion transformer with temporally reversible motion tokenizations; (2) We develop Adv-BMT for realistic and diverse safety-critical traffic simulations; (3) We leverage Adv-BMT in a closed-loop setting to dynamically create challenging environments for reinforcement learning agents.
@@ 2 Related Work
Driving Motion Prediction.
The task of motion prediction focuses on forecasting the future trajectories of traffic participants conditioned on their initial map context and agent states. Recent work uses transformer models to autoregressive sequence modeling. A set of work utilizes discretized motion tokenizations to perform next token predictions: Trajeglish philion2024trajeglishtrafficmodelingnexttoken discretizes motion representations using a k-disk-based tokenization scheme to represent position and angle differences for relative movements. MTR++ shi2024mtrmultiagentmotionprediction directly represents motion on continuous space in Gaussian mixture distributions. MotionLM seff2023motionlmmultiagentmotionforecasting models trajectory deltas. SMART builds a k-disk–clustered vocabulary of motion tokens, containing coordinates, heading, and shape. BehaviorGPT zhou2024behaviorgptsmartagentsimulation performs next-patch predictions with future motion chunks, instead of single-step predictions. The BMT model constructs two sets of motion tokens based on inter-frame accelerations, enabling both forward prediction of future motions and inverse reconstruction of past motions. Another set of works Jiang_2023_CVPR ; zhong2022guidedconditionaldiffusioncontrollable ; zhong2023languageguidedtrafficsimulationscenelevel ; 10610324 ; 11127600 ; NEURIPS2023_d95cb79a ; Xu_Petiushko_Zhao_Li_2025 ; 10801623 leverage diffusion models for motion predictions.
Safety-Critical Traffic Scenario Generation.
@@
To generate collision traffic, a classic line of previous works use a two-staged method: first uses a traffic prior to generate realistic agent trajectories, then use collision-sensitive objectives for trajectory optimization or candidate selections. STRIVE rempe2022generatingusefulaccidentpronedriving models traffic by a variational autoencoders but requires computationally expensive per-scenario optimizations. CAT zhang2023catclosedloopadversarialtraining leverages transformer motion decoder gu2021densetntendtoendtrajectoryprediction and selects the collision trajectories and simulates in MetaDrive li2022metadrivecomposingdiversedriving for closed-loop adversarial environments. SEAL stoler2025sealsafeautonomousdriving uses a skill-based adversarial policy with collision-related objectives. SafeSim chang2024safesimsafetycriticalclosedlooptraffic proposes a diffusion model with a test-time collision-sensitive guidance loss to control the collision type and adversarial agent selections. Another set of work such as ransiek2024goosegoalconditionedreinforcementlearning use a reinforcement learning (RL) based approach, which parametrizes trajectories and goal constraints to generate safety-critical interactions. AdvSim wang2023advsimgeneratingsafetycriticalscenarios directly perturbs actor trajectories using a kinematic model and optimizes via a black-box adversarial loss. Another line of recent work such as ding2020multimodalsafetycriticalscenariosgeneration uses a conditional normalizing flow to model the distribution of real-world safety-critical trajectories. CrashAgent li2025crashagentcrashscenariogeneration and LCTGen tan2023languageconditionedtrafficgeneration leverage free-form texts as inputs and extract embeddings to parameterize scene initializations and agent driving directions. Different from previous approaches, Adv-BMT generates collision scenarios in three steps: first samples a collision state, then conduct reverse predictions, finally forcast traffic agents accordingly.
3 Method
Classical motion prediction models forecast future trajectories based on the current states of traffic agents. Building on this foundation, we propose the reverse motion prediction problem. To address both tasks within a unified framework, we introduce the Bidirectional Motion Transformer (BMT) model, which is able to perform both tasks. Finally, we present the Adv-BMT framework and describe how it leverages BMT model as the core to generate realistic and diverse safety-critical interactions.
Figure 2: Bidirectional predictions on the ego agent (red). BMT supports predictions for future motions (left) and historical motions (right) jointly for all prediction agents.
3.1 Bidirectional Motion Prediction
@@
We first introduce the bidirectional motion prediction task, shown in Fig. 2. Consider a traffic scenario with at most $N$ agents and a prediction horizon of $T$ steps. The trajectory of agent $i$ is represented as $\boldsymbol{\tau}^{i}=\{\tau^{i}_{0},\tau^{i}_{1},\dots,\tau^{i}_{T}\}$, where each state $\tau^{i}_{t}\in\mathbb{R}^{d}$ encodes its position, velocity, and heading at time $t$. We introduce a prediction direction indicator $D\in\{\text{Forward},\text{Reverse}\}$, to specify whether the model predicts future or past motion over the horizon.
For each predicted traffic agent, we construct a sequence of motion tokens $\mathbf{Z}^{i}=\{z^{i}_{1},\dots,z^{i}_{T}\}$, by applying the motion tokenization function $\phi(\cdot)$ between consecutive states. In the forward setting, the tokens are derived in chronological order, from $\tau^{i}_{0}$ to $\tau^{i}_{T}$. In the reverse setting, the temporal order of the trajectory is inverted, and the tokens are generated by applying $\phi(\cdot)$ backward from the current state toward the initial state. This formulation yields a bidirectional token sequence that enables BMT to model both forward and reverse motion dynamics in a unified token space. The tokenization function $\phi(\cdot)$ is further detailed in Section 3.2.
3.2 Bidirectional Motion Transformer (BMT)
Token Space.
@@ 3.2 Bidirectional Motion Transformer (BMT)
Motion Token Reconstruction.
A key step in BMT is to reconstruct continuous trajectories from discrete motion tokens, bridging between physical motion and the token vocabulary. Each token encodes a low-level control command defined by an acceleration $a$ and a yaw rate $\omega$. Starting from the current agent state $\tau=(x,y,\theta,v)$, the token is applied over a time interval $\Delta t$ to update the agent’s position, heading, and velocity. BMT adapts midpoint integration to propagate the state, which provides a stable and realistic approximation of vehicle dynamics. Intuitively, the speed and heading are updated using their midpoint values, ensuring that the reconstructed trajectory remains smooth and physically consistent. This reconstruction procedure allows BMT to decode token sequences into continuous motion trajectories in both forward and reverse prediction settings.
Model Architecture.
@@
The BMT architecture overview is shown in Fig. 3. BMT has a scene encoding component used to obtain embeddings for scenario contexts with separate embeddings for map polylines, traffic lights, and agent initial states. Then, we use Fourier-encoded edge features tancik2020fourierfeaturesletnetworks to represent the spatial and directional information between these encoded entities, which are then passed to the transformer encoder with self-attention layers for the relational embeddings.
Figure 3: BMT architecture. BMT consists of a scene encoder and a GPT-style motion decoder. It employs two sets of motion tokens for forward and reverse predictions to generate the next-step token for each agent. All predictions are conditioned only on the map information and the one-step current state of all predicted agents.
@@
The prediction decoder predicts subsequent motion tokens in an autoregressive manner, with only initial agent information for the first frame, along with the scene embedding obtained from the Scene Encoder. The motion decoder incorporates self-attention over the initial agent token embeddings, and three relation computations separately: agent-to-agent (a2a), agent-to-time (a2t), and agent-to-scene (a2s), with each relation embedding then passed to its cross-attention layers. The output agent embeddings are concatenated and repeated a number of times. The output agent motion embeddings are mapped to the vocabulary of discretized motion tokens through MLPs. More details can be found in the Appendix.
Training.
BMT is trained to learn a policy that reproduces the distribution of real-world driving behaviors. At each step, the model predicts a discrete motion token for every agent based on its past tokens, current state, and the scene context. To align these predictions with the ground-truth behaviors in the dataset, we minimize a cross-entropy loss between the predicted and observed token distributions:
@@
$\mathcal{L}_{\text{train}}=-\;\mathbb{E}_{\mathcal{D}}\left[\sum_{t=1}^{T}\sum_{i=1}^{N}\log\pi_{\theta}\left(z^{i}_{t}\mid\mathbf{z}^{i}_{1:t-1},\mathcal{M}\right)\right],$ (1)
where $\pi_{\theta}$ is the token prediction policy parameterized by $\theta$. Intuitively, this objective encourages the model to assign high probability to motion tokens that correspond to real trajectories, thereby capturing the joint distribution of multi-agent actions in traffic.
During inference, BMT generates motion tokens autoregressively, sampling one token at a time conditioned on its previous predictions. We apply nucleus (top-$p$) sampling to promote behavioral diversity while remaining faithful to the learned distribution. To mitigate exposure bias, the model rolls out on its own sampled tokens rather than the ground-truth sequence, ensuring consistency between training and inference.
3.3 Adv-BMT for Safety-critical Interaction
The BMT model serves as the core that captures the distribution of realistic multi-agent motions in Adv-BMT framework. Building upon this foundation, we develop Adv-BMT, an adversarial scenario generation framework designed to generate realistic and safety-critical traffic interactions. The overview of Adv-BMT framework is illustrated in Fig.1. From the input scenario, Adv-BMT first samples diverse collision states between ego agent and a new adversarial agent (ADV), then reconstructs complete trajectories via reverse motion predictions. A rule-based reject sampling mechanism is used for selecting collision initializations.
Figure 4: Diverse collision headings generated by Adv-BMT from the same driving log input, visualized from a third-person perspective.
Figure 5: Diverse collision timings generated by Adv-BMT from the same driving log input, visualized in a bird’s-eye view.
Diverse Adversarial Initialization.